OpenStudy (anonymous):

Find two Arithmetic means between -5 and 4

OpenStudy (anonymous):

The arithmetic mean of these two is (-5 + 4)/2 =

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

Definition of arithmetic mean of n numbers is the sum of those numbers then divided by n.

OpenStudy (anonymous):

then add the mean to -5? Because it asks for two arithmetic means in between.

OpenStudy (anonymous):

@Seabass and @douglaswinslowcooper, to find two arithmetic means we can first find the distance between the points -5 and 4. distance = |-5 - 4| = |-9| = 9. Then we can divide by 3 to find the distance from -5 to the first arithmetic mean. 9/3 = 3. First arithmetic mean: -5 + 3 = -2 Second arithmetic mean -2 + 3 = 1 The arithmetic means are -2 and 1.

OpenStudy (anonymous):

awesome, that exactly what i was thinking

OpenStudy (anonymous):

This might make more sense if we think of cutting the interval into thirds instead of halves.

OpenStudy (anonymous):

if you elaborated anymore if would probaly overheat and my brain would explode so we'll leave it at that. Thanks alot. I just needed a clarification if anything.